OpenStudy (anonymous):

a ball is thrown vertically upward w/a velocity of 20m/s find a.time of ascent b.total time in air c.max height reached d.up when it reaches the point of release

OpenStudy (anonymous):

Try using one of the 4 Newtonian equations of motion, knowing the initial velocity, and acceleration due to gravity (this will turn into a quadratic for part a): \[y = y_0 + v_0*t + 1/2*a*t^2\]

OpenStudy (anonymous):

well in this case its a deceleration against gravity, also we need to find time, So better equation to use for point a is v=u-gt, v being the final velocity which is zero, u being the initial velocity (given) and g being acceleration due to gravity, find t. Hope this helped

OpenStudy (anonymous):

Mashy, while you are on the correct path, I believe solving your equation will only give you approximately half the time it is in the air, since v also equals 0 at the top of the vertical trajectory, and only reaches a final velocity of zero after it has impacted and rested. As a result, the instantaneous velocity at y = 0 at the end of its flight will not be zero.

OpenStudy (anonymous):

no no.. Time of ascend means the time until which the ball stops going upwards and will start to come down, So before it starts falling, it has reached the velocity of zero, so by solving the equation the time which you get is exactly the time it took the ball to reach its maximum height, which is nthing but time of ascend!

OpenStudy (anonymous):

why not just use vf = vi + at and solve for t given that a is (-9.8) because it is an acceleration opposite to the initial velocity, therefore we can use 0 = 20 + (-9.8)t and solve for t which after solving is 2.04 seconds

OpenStudy (anonymous):

Thats exactly what I said :D

OpenStudy (anonymous):

But lets not give answers! Only tips and hints to solve!

OpenStudy (anonymous):

Total time in the air is 4.08, double the time required for it to stop. The max height can be found using \[v _{f}^{2} = v_{i}^{2} + 2a \Delta d\] solving for d of course assuming vf to be 0 and vi to be 20 and a to be (-9.8)

OpenStudy (anonymous):

We arent supposed to solve them :S

OpenStudy (anonymous):

my bad then lol

OpenStudy (anonymous):

yea, it could be their homework problems, in which case we are not helping but spoiling, so only give clues to help them understand better not so much as to blindly copy it :D

OpenStudy (anonymous):

a = 2.0408163265306122448979591836735 s b = 4.0816326530612244897959183673469 s c = 20.408163265306122448979591836735 m what do you mean by d? -20 m/s down is the velocity 0 m is the displacement 40.816326530612244897959183673469 m is total distance traveled and I can keep going but i think d question is mistyped

OpenStudy (anonymous):

the duke - what's with all the decimal places?????? physics is not math; numbers (measurements) aren't exact!

OpenStudy (anonymous):

You can easily calculate these out......Just use Neuton's law of motion. \[v ^{2}=u ^{2}+2fs\]where v=0, You know f=(-9.8m/s). You can know the value of s(the height)......And now measure the time easily...... You have these formulas-\[v=ut+(1\div2)ft ^{2}\] go ahead.......

OpenStudy (anonymous):

the complete solution with explanations is attached below in a pdf file.

OpenStudy (anonymous):

@psd 007..... man wat a complex method .. we know that V=U+at at highest point V=0 initial speed U=20 a= -9.8 so time of accent= 20/9.8= 2.04... One line answer . time of accent = time of decent => total time =2*time of accent =4.08 Once U Have the times its olways very easy to use newtonian equations to calculate rest everything else . Hope It'll be Easy For you @KingKos to find the soltion now :)

OpenStudy (anonymous):

@manjotpalsingh1993: You are right, your method is far shorter than mine. Good job. I think that my method reduces to yours, so really they are the same, yours is just faster. Thanks for this. :)